# FedMerge: Federated Personalization via Model Merging

### v2

@@ 4.1 Classical FL Settings
Table 2: Comparison with Single-model Methods on Client-resource Constrained Scenarios. We use ResNet-9 for all methods.
Non-IID Local FedAvg FedAvg+ pFedMoE FedMerge 5 10 15 20 25 30 Personal 38.76 30.29 53.54 54.28 37.88 46.62 51.18 53.52 53.99 54.38 Cluster 23.13 29.20 46.53 49.03 51.69 51.38 53.27 51.16 51.69 51.96 Domain 54.97 60.19 60.39 63.59 62.12 64.64 65.80 65.32 65.26 64.96
Table 3: Numerical Results for Multi-Model Methods by Varying global model Numbers. We use ResNet-9 for all methods.
Non-IID Method Global Model Num
5 10 15 20 25 30
Personal IFCA 25.68 26.45 28.86 27.35 29.68 30.80
FedEM 32.50 33.67 34.59 37.08 38.12 39.48
FedMerge 37.88 46.62 51.18 53.52 53.99 54.38
Cluster IFCA 42.64 41.06 44.97 40.34 37.53 38.04
FedEM 48.47 48.34 46.35 46.16 46.51 46.65
FedMerge 51.69 51.38 53.27 51.16 51.69 51.96
Domain FedEM 61.77 61.33 64.63 61.23 62.93 63.14
FedMerge 62.12 64.64 65.80 65.32 65.26 64.96
Non-IID Settings FedMerge adapts to different Non-IID settings by adjusting the merging weights adaptively. To validate this adaptability, we consider three data distributions using the CIFAR-100 (Krizhevsky et al., 2009) and PACS (Li et al., 2017) datasets: Cluster Non-IID: We split the original CIFAR-100 dataset into 5 clusters based on label categories. Each cluster is evenly assigned to a subgroup of clients out of 50 clients. The ground-truth clustering is {{1, …, 6}, {7, …, 11}, {12, …, 19}, {20, …, 32]}, {33, …, 50}}. Personalize Non-IID: A commonly used Non-IID setting proposed by (Hsu et al., 2019), where the CIFAR-100 dataset is partitioned using a Dirichlet distribution into 50 clients. We set the concentration parameter $\alpha=0.1$ for high Non-IIDness. Domain Non-IID: We use the PACS (Li et al., 2017) dataset, assigning each client to one domain, resulting in 4 clients.
FedMerge Settings The learning rate of the global models $\Theta$ is set to 1 for all clients, meaning that at each round, the global model is replaced with the local update model. For updating the merging weights, we use a fixed-magnitude update strategy to ensure smooth parameter updates. Specifically, at each round, we perform a grid search on the gradient direction with small steps and stop when the update of the merging weight reaches 0.01.
@@
System Settings For all methods, we split the dataset into training, validation, and test sets, selecting the checkpoint with the minimum validation error. The number of communication rounds is set to 500 to ensure convergence. Each method is executed three times, and the reported results are the average of these runs.
4.2 Numerical Comparison for Multi-Model Methods
As shown in Table 1, for IFCA, FedEM, and FedMerge, we vary the number of global models chosen from the set $\{5,10,15,20,25,30\}$, using ResNet-9 (He et al., 2016) as the basic model. Note that we do not run IFCA on Domain Non-IID since having only 4 clients is too few for Cluster FL methods.
Table 3 suggests the following: 1) Multi-Model methods can boost their performance by using a large number of models. For example, on Personal Non-IID, the performance continues to improve for all methods as the number of models increases. On Cluster Non-IID and Domain Non-IID, FedMerge achieves optimal performance with $15$ global models. 2) FedMerge consistently surpasses FedEM and IFCA across various global model numbers, demonstrating the effectiveness of FedMerge compared to other Multi-Model FL methods.
4.3 Numerical Comparison for Single-Model Methods in Client-Resource-Constrained Scenarios
For Single-model methods (Local, FedAvg, FedAvg+, pFedMoE), performance can be improved by using larger models. However, increasing model scale significantly raises client resource consumption, which is a critical constraint in FL. To ensure a fair comparison, we first analyze the performance of Single-model methods under client-resource-constrained scenarios. Then, we analyze all methods under unlimited resource conditions.
FedMerge can cheaply boost performance by sharing a large number of models on the server without requiring much client resources. In Table 2, we compare FedMerge with Single-model methods under constrained client resources. For all baselines, each client communicates one ResNet-9 model with the server. All methods, except pFedMoE, optimize one model on each client, while pFedMoE optimizes two models. We assume server resources are unlimited, allowing FedMerge to boost performance by using a large number of global models. As shown in Table 2, on Cluster and Domain Non-IID, FedMerge consistently outperforms all baselines across all numbers of global models. On Personal Non-IID, although FedMerge performs poorly with a small number of global models (5), its performance consistently increases as the number of global models grows, eventually surpassing pFedMoE with 30 global models.
Figure 3: Server Parameter Usage vs. Accuracy. The x-axis represents multiples of ResNet-9’s parameter count. In FedMerge, we fix the architecture to ResNet-9 and vary the number of global models from [5, 10, 15, 20, 25, 30].
4.4 Server Parameter Usage vs. Performance
We analyze the parameter efficiency of both single-model and multi-model methods under the assumption of unlimited server resources. All methods follow the parameter scaling configurations outlined in Table 1.
Fig. 3 presents server parameter usage versus accuracy for Personal Non-IID and Cluster Non-IID settings. A comprehensive comparison across all Non-IID scenarios is provided in the Appendix. The results indicate the following:
FedMerge trades off higher server-side resource consumption for improved client-side efficiency. In Fig. 3, parameter-efficient methods, which achieve high performance with minimal server parameter usage, appear in the top-left corner. Multi-Model methods are less server resource-efficient compared to Single-model methods, as they utilize more parameters on the server while achieving comparable performance. However, FedMerge and IFCA require transmitting and optimizing only a single ResNet-9 per client. Despite FedMerge utilizing substantial server resources in Fig. 3, it remains client-efficient and achieves the highest performance among all baselines. Notably, pFedMoE assigns two models and an additional routing network to each client, whereas FedMerge allocates only a single model. As a result, FedMerge outperforms pFedMoE in Personal Non-IID with 30 global models and surpasses pFedMoE across all settings in Cluster Non-IID.
@@ 4.3 Numerical Comparison for Single-Model Methods in Client-Resource-Constrained Scenarios
4.5 Client Parameter Usage vs. Performance
@@
From Fig. 3, the performance of single-model methods degrades as model scale increases. This is because both Personal and Cluster Non-IID settings involve 50 clients, where the large number of clients with sparse client data limits performance gains from scaling up the model. In contrast, Domain Non-IID is more conducive to model scaling, as it consists of only four clients with abundant local data.
@@
4.6 Ablation Study
FedMerge adapts to Non-IID settings with dynamic merging weights: To understand the behavior of FedMerge under different Non-IID settings, we visualized the heatmaps of merging weight matrices under different communication rounds for Personal Non-IID and Cluster Non-IID in Fig. 5. Each row represents the normalized merging weights for a single client. Initially, the merging weight matrix is uniformly initialized, and as the epoch progresses, it begins to exhibit specific distribution patterns. We separate the ground-truth clusters of Cluster Non-IID with white dashed lines. In Cluster Non-IID, rows of merging weights are similar within each ground-truth cluster and distinct between different clusters. This indicates that FedMerge automatically allows clients with similar data distributions to prefer similar subgroups of global models, achieving a soft clustering. In contrast, for Personal Non-IID, rows of merging weights are relatively distinct from one another, and the merging weight matrix does not exhibit a clear structure.
Table 4: Ablation study on merging weights. Fixed indicates that each client randomly select some global models. The corresponding merging weight is uniformly initialized and frozen during training. ResNet-9 is used for all methods.
Non-IID FedAvg FedMerge (15 global models)
Fixed (1/15) Fixed (5/15) Fixed (10/15) Dynamic
Personal 29.20 40.94 44.7 45.48 51.18
Cluster 30.29 32.55 32.52 34.06 53.27
Table 5: Numerical results on foundation models with multi-language datasets.
Dataset Local FedIT FedLoRA FedMerge
XNLI 0.698 0.770 0.769 0.783
XQUAD 0.561 0.559 0.556 0.566
This irregularity aligns with the nature of Personal Non-IID, where each client has a unique data distribution.
Effect of different weight merging strategies: We further analyze different weight merging strategies in Table 4. Specifically, each client randomly select a subset of models out of 15 global models. The corresponding merging weights are randomly initialized and frozen. As observed, The performance of Fixed strategy decreases compared with dynamic merging, and Cluster FL decreases more than Personal Non-IID because random selection conflicts with the structural information of the cluster.
4.7 Foundation-model FL Settings and Results
In federated learning (FL) with foundation models, we validate FedMerge with a group of tasks, where each client is assigned a distinct task. While these clients share commonalities, they also have distinct features that reflect their specific styles. Specifically, we use datasets from multi-language tasks, XNLI (Conneau et al., 2018) and XQUAD (Artetxe et al., 2019). Each client is assigned a single language, and the content of the client data is identical across languages, differing only in the language used. For XNLI, we randomly select 300 samples for training and 200 samples for evaluation for each client. For XQUAD, we skip samples that exceed the maximum token length, randomly select 600 samples for training, and use the remaining samples for evaluation. The batch size is set to 8, and the maximum token length is 512.
We compare FedMerge with three baselines: Local, FedIT (Zhang et al., 2024), and FedLoRA (Yi et al., 2023). The ”Local” baseline refers to models trained exclusively on local data without any communication between clients or the server. We adapt Alpaca-LoRA [^1] https://github.com/tloen/alpaca-lora as the base model, initialized with LLaMA3-8B (Dubey et al., 2024). The rank of LoRA is set to 8. For FedMerge, we use 5 global adapters. ROGUE-1 is used as the evaluation metric. The number of communication rounds is set to 20, and the checkpoint with the minimum validation error is selected.
As shown in Table 5, on XNLI, sharing information among clients significantly enhances performance (e.g., FedAvg vs. Local), indicating that knowledge from common language clients (e.g., English) helps the model understand rare language clients. On XQUAD, sharing information among clients using a single model does not enhance performance (FedAvg vs. Local). FedMerge consistently outperforms the compared methods on both multi-language datasets, demonstrating the effectiveness of FedMerge in FL with foundation models for multi-language tasks.
4.8 Conclusion
In this paper, we propose an end-to-end Model Merging framework for Federated Learning, where merging weights and global models are optimized jointly from scratch. The server manages all global models and merging weights, customizing personalized models for each client by weight-averaging the global models. FedMerge can boost performance with a large model soup on the server, while still being client-resource friendly by only communicate and optimize a single model for each client.
FedMerge offers a promising direction for Federated Foundation Models: A large server knowledge pool can match foundation model scales, while each client selects and combines a small subset of parameters without transmitting the full model.
Appendix A More Discussion on Parameter Usage vs. Accuracy
We analyze three different parameter usage standards: global parameter scale on the server (Server), communicated parameter scale for each client (Communicate), and stored parameter scale for each client (Client). We adopt the parameter scale settings in Table 1.
As a federated learning (FL) method with multiple models, FedMerge requires storing and optimizing a large number of parameters on the server, which limits its parameter efficiency in terms of global models compared to single-model FL methods. As shown in the first row of Fig. 7, FedMerge is less parameter-efficient on the server side than single-model methods. However, in FL, client resources are typically more valuable than server resources. FedMerge can utilize a large number of global models without significantly increasing client resource usage.
To illustrate this, we plot communicated parameter usage vs. accuracy and client-optimized parameter usage vs. accuracy in Fig. 7. From the second and third rows in Fig. 7, FedMerge and IFCA exhibit straight lines at the top-left corner. This is because these methods use only a single model for each client, enabling performance improvement without increasing client resource requirements.
Figure 7: Parameter usage vs. accuracy across all Non-IID settings and parameter standards. ”Server” refers to parameter usage on the server, ”Communicate” denotes the number of parameters communicated by each client in a single round, and ”Client” represents parameter usage for each client. Points closer to the top-left corner indicate higher performance with a smaller parameter scale, which is more efficient.
Appendix B More Discussion on Merging Weight Matrix
@@
We provide additional visualizations of the merging weight matrix in Fig. 8. Specifically, in Fig. 8(b), the evolution speed of different clusters varies. Smaller clusters with fewer clients evolve faster and demonstrate distinct patterns within 200 epochs, as observed in the first three clusters with small client IDs. However, for the largest cluster (clients with IDs [32, …, 50]), the merging weights evolve more slowly, even after 500 epochs. This is because the largest cluster has the highest number of clients (19 clients) and requires more global models, resulting in a more uniform merging weight distribution.

### v3

@@ Introduction
Foundation Model FL (LoRA-based)
Multi-Model Description Rank-8×2 Rank-8×4 Rank-8×6 Rank-8×8 Rank-8×10 Rank-8×12
Parameter Scale 2 4 6 8 10 12
Single-Model Description Rank-8 Rank-16 Rank-24 Rank-32
Parameter Scale 1 2 3 4
Table 1: Parameter scale settings for classic FL (top) and foundation model FL (bottom). The scale is measured relative to Rank-8 for LoRA modules and to ResNet-9 for ResNet architectures.
Technical Implementations
Constraint on Merging Weights Values The original objective of FedMerge in Eq (4) has no constraints on merging weights, allowing the merging weight values to be negative. However, neural network models typically require regularized parameters. To address this, we apply a softmax function to the merging weights to obtain normalized weights. This introduces slight modifications to the updates in Eq (8) and Eq (9). Detailed derivations can be found in Appendix D.
Clarification of Merging Weights Gradient In Eq (9), the update of merging weights involves the inner product of model parameters and gradients. For large models, the effective update directions can be overwhelmed by the inner products of a large number of parameters. To mitigate this issue, we simply use the parameters from the last few layers of the model for inner product—specifically, the classification head in convolutional neural networks. This is because higher-level layers capture more semantic information (Yosinski et al. 2014), and their parameters or gradients are correlated with task heterogeneity across clients.
Server-side Computational Complexity
FedMerge is as computational efficient on the server side as standard baselines Server-side model assembly may raise concerns about computational overhead. We compare the server-side complexity of FedMerge with widely adopted baselines. As in Eq (4), Eq (8) and Eq (9), FedMerge performs weight averaging $d$ times over $m$ models or computes $md$ inner products, resulting in a complexity of $\mathcal{O}(md)$. Cluster FL (Ghosh et al. 2020; Long et al. 2023) and Multi-model FL (Marfoq et al. 2021; Bhuyan and Moharir 2022)—both widely accepted in the literature—also require performing weight averaging $d$ times over $m$ models on the server side, resulting in a complexity of $\mathcal{O}(md)$. Thus, FedMerge does not introduce additional computational complexity on the server side compared to these standard approaches.
Experiments
Classical FL Settings
Baselines FedMerge relates to Cluster FL, Multi-model FL, and MoE-like FL. We compare it with the following representative methods: 1) Single-model methods share a single global model, including Local (training local models separately on each client), FedAvg (McMahan et al. 2017), FedAvg+ (FedAvg with local fine-tuning), and pFedMoE (Yi et al. 2024). Although named “MoE,” pFedMoE uses only one global model. 2) Multi-model methods maintain multiple global models, such as IFCA (Ghosh et al. 2020) (Cluster FL) and FedEM (Marfoq et al. 2021) (MoE-like FL).
Parameter Scale To ensure fair comparison, we control the parameter scale for each method. Multi-model methods (IFCA, FedEM, FedMerge) use ResNet-9 with varying numbers of global models from {5, 10, 15, 20, 25, 30}. Single-model methods (Local, FedAvg, FedAvg+, pFedMoE) adopt different ResNet variants: ResNet-9, ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152, whose parameter scales relative to ResNet-9 are {1, 2.27, 4.32, 4.82, 8.73, 11.95}. Full settings are shown in Table 1.
FedMerge Settings Global models are randomly initialized and trained from scratch. Merging weights are initialized uniformly to reflect equal initial contributions from global models.
@@
System Settings These are the universal settings for all methods across all non-IID settings. Each dataset is split into training, validation, and test sets, with the best checkpoint selected based on validation error. We use a constant learning rate of $0.01$, local training for $2$ epochs, and $500$ communication rounds. Each method is executed three times, and the reported results are the average of these runs.
Parameter Usage vs. Performance
Figure 3: Server parameter usage vs. accuracy. The x-axis represents multiples of ResNet-9’s parameter count. For each method, the variation of server parameters follows Table 1. The highest and second-highest accuracy points are marked with black and gray respectively.
@@
FedMerge adapts to different Non-IID settings by adjusting the merging weights adaptively. To validate this adaptability, we consider two data distributions of different non-IID structures:
Non-IID Settings We build two non-IID settings using the CIFAR-100 (Krizhevsky et al. 2009) dataset: Cluster Non-IID: We split the original CIFAR-100 dataset into $5$ clusters based on label categories. Each cluster is evenly assigned to a subgroup of clients out of 50 clients. The ground-truth clustering is {{1, …, 6}, {7, …, 11}, {12, …, 19}, {20, …, 32}}, {33, …, 50}}. Personal Non-IID: A commonly used Non-IID setting proposed by (Hsu et al. 2019), where the CIFAR-100 dataset is partitioned using a Dirichlet distribution into 50 clients. We set the concentration parameter $\alpha=0.1$ for high Non-IIDness.
Following Table 1, we vary Single-model methods by model architecture and vary Multi-model methods by number of models, which results in various server parameter usage. As shown in Fig. 3, we plot the server parameter usage vs. accuracy on Personal Non-IID and Cluster Non-IID settings. The results suggest the following:
Multi-model FL methods provide an alternative way to improve performance beyond scaling model architectures. In Fig. 3(a), using larger ResNet architectures for Single-model methods leads to a decrease in performance, while using more ResNet-9 models for Multi-model methods can boost the performance (with the largest architecture considered being ResNet-512 for Single-model methods, and up to 30 models used for Multi-model methods). This is because both Personal and Cluster Non-IID settings involve 50 clients with sparse client data and high non-IID data partition settings, which limits performance gains from scaling up the model. In contrast, Multi-model methods can enhance the base model’s ability to handle complex non-IID distributions by increasing the number of models. FedMerge consistently surpasses FedEM and IFCA across various global model numbers, demonstrating the effectiveness of FedMerge compared to other Multi-Model FL methods.
FedMerge boosts performance by shifting computation to the server, leveraging the server’s abundant resources to accommodate the limited client-side capacity. In Fig. 3, FedMerge manages a number of ResNet-9 model on the server while transmitting and optimizing a single model per client. Notably, on Personal Non-IID setting, FedMerge outperforms pFedMoE under the Personal Non-IID setting $(54.38>54.34)$ with only one ResNet-9 per client, while pFedMoE needs two ResNet-18 models and additional router weights per client to achieve optimal performance. On Cluster Non-IID, FedMerge consistently surpasses other baselines and achieves optimal performance with $15$ global models.
We also provide Non-IID settings with PACS (Li et al. 2017) dataset, where using larger ResNet architectures can boost performance. Additional results are in Appendix A.
Ablation Study
FedMerge adapts to Non-IID settings with dynamic merging weights: To understand the behavior of FedMerge under different Non-IID settings, we visualized the heatmaps of merging weight matrices in Fig. 4. Each row represents the normalized merging weights for a single client. Initially, the merging weight matrix is uniformly initialized, and as the epoch progresses, it begins to exhibit specific distribution patterns. We separate the ground-truth clusters in Cluster Non-IID with white dashed lines. In Cluster Non-IID, rows of merging weights are similar within each ground-truth cluster and distinct between different clusters. This indicates that FedMerge automatically allows clients with similar data distributions to prefer similar subgroups of global models, achieving a soft clustering. In contrast, for Personal Non-IID, rows of merging weights are relatively distinct from one another, and the merging weight matrix does not exhibit a regular structure.
Non-IID FedAvg FedMerge (15 global models)
Fixed (1/15) Fixed (5/15) Fixed (10/15) Dynamic
Personal 29.20 40.94 44.70 45.48 51.18
Cluster 30.29 32.55 32.52 34.06 53.27
Table 2: Ablation study on merging weights. Fixed indicates that each client randomly selects some global models, and the corresponding merging weight is uniformly initialized and frozen during training.
Method Paraphrase Entailment Structure to text Text formatting Linguistic acceptability Word disambiguation Coreference Question classification Average
FedEM 73.40 74.30 69.32 87.93 78.80 61.31 66.82 78.33 73.77
FedMerge 75.49 77.02 68.54 86.97 80.64 59.70 70.61 80.03 74.86
Table 3: ROUGE-1 scores (%) of MoE-like method (FedEM) and model merging (FedMerge) in task-heterogeneous instruction following. Results are reported across 8 instruction-following tasks from Flan, with the average shown in the last column.
Effect of different weight merging strategies: We further analyze different weight merging strategies in Table 2. As observed, The performance of Fixed strategy decreases compared with dynamic merging, and Cluster FL decreases more than Personal Non-IID because random selection conflicts with the structural information of the cluster.
Federated Finetuning of Large Language Models
Baselines We use the same baselines as in classic FL settings, except that FedAvg is replaced with FedIT (Zhang et al. 2024) and FedAvg+ is replaced with FedLoRA (Yi et al. 2023). For pFedMoE and FedEM, we build personalized routers for the LoRA modules. Details can be found in Appendix B.
Datasets Several baselines (Ye et al. 2024b; Kuang et al. 2024; Ye et al. 2024a) exist for federated foundation models. However, these models are primarily designed for instruction fine-tuning, and their goal is to learn a global model that follows general instructions rather than to personalize the LLM’s behavior for each client. In this paper, we consider two datasets for personalized LLM: Task-heterogeneous instruction following: Following (Yang et al. 2024b), we randomly select $8$ NLP tasks from Flan (Wei et al. 2021). One task per client, resulting in $8$ total clients. Each task randomly selects $300$ training samples and $200$ test samples. ROUGE-1 is used as the evaluation metric. Persona-aligned instruction following: We also provide a more realistic scenario with RoleBench (Wang et al. 2023), where each client takes on a specific role and has a Q&A dataset with clear role characteristics. Details can be found in Appendix B.
System Settings We adapt Alpaca-LoRA [^1] https://github.com/tloen/alpaca-lora as the base model, initialized with LLaMA3-8B (Dubey et al. 2024). The rank of LoRA is set to $8$. We employ a linear learning rate decay schedule from 5e55e{-}5 to 5e65e{-}6 over the training process. The number of communication rounds is set to $50$ to ensure convergence. For each round, the local training is conducted with $20$ steps, and the batch size is set to $4$.
Parameter Scale Similar to classic FL settings, we vary the number of LoRA modules for Multi-model methods and vary the rank of LoRA modules for Single-model methods. Details are shown in Table 1. For cluster FL methods like IFCA, the number of global models is limited to at most $8$, equal to the number of clients. All experiments are conducted on a single NVIDIA A40 GPU.
Figure 5: Server parameter usage vs. accuracy on the federated Flan dataset. The values on the x-axis are shown as multiples of the parameter count of a Rank-8 LoRA. The highest and second-highest accuracy points are marked with black and gray respectively.
Following Table 1, we vary Single-model methods by the LoRA rank and Multi-model methods by the number of LoRA modules. The server parameter usage vs. accuracy for the federated Flan dataset is shown in Fig. 5. FedMerge demonstrates better instruction-following abilities on the federated Flan dataset by sharing up to $12$ LoRA modules globally while only transmitting one LoRA module of rank $8$ per client. Results for the federated RoleBench dataset can be found in Appendix B.
Model Merging (FedMerge) provides a valuable alternative to MoE-like methods (FedEM) for FL with foundation models while being more client-resource friendly. Unlike classic settings, MoE-like methods are widely used in foundation models. A typical setup for a Mixture-of-Experts (MoE) foundation model in a federated learning scenario is that each client hosts several experts and a local router. These experts are sent to the server and aggregated separately to form a pool of global experts. While being simple in optimization strategy, it is not as client-resource-efficient as FedMerge. A detailed comparison between FedEM and FedMerge is presented in Table 3. We select the best performance for FedEM and FedMerge within the rank settings in Table 1, which is $12$ shared LoRA modules for FedEM and $8$ for FedMerge. From Table 3, FedMerge and FedEM outperform each other on different tasks, while FedMerge is better in average performance across clients.
Conclusion
In this paper, we propose an end-to-end Model Merging framework for Federated Learning, where merging weights and global models are optimized jointly from scratch. The server manages all global models and merging weights, customizing personalized models for each client by weight-averaging the global models. FedMerge can boost performance with a large model soup on the server, while still being client-resource friendly by only communicating and optimizing a single model for each client. FedMerge offers a promising direction for Federated Foundation Models: a large server knowledge pool can match foundation model scales, while each client selects and combines a small subset of parameters without transmitting the full model.
Appendix A More Experiments on Classic FL
Client Parameter Usage vs. Performance
@@ Appendix A More Experiments on Classic FL
From Fig. 3, the performance of single-model methods degrades as model scale increases. This is because both Personal and Cluster Non-IID settings involve 50 clients, where the large number of clients with sparse client data limits performance gains from scaling up the model. Therefore, we build Non-IID setting with less heterogeneity. Each client is assigned sufficient number of samples to ensure scaling up model architectures more effective:
Domain Non-IID: We use the PACS (Li et al. 2017) dataset, assigning each client to one domain, resulting in 4 clients. Note that we do not run IFCA on Domain Non-IID since having only 4 clients is too few for Cluster FL methods.
@@
More Discussion on Merging Weight Matrix
@@
We provide additional visualizations of the merging weight matrix in Fig. 7. Specifically, in Fig. 7(b), the evolution speed of different clusters varies. Smaller clusters with fewer clients evolve faster and demonstrate distinct patterns within 200 epochs, as observed in the first three clusters with small client IDs. However, for the largest cluster (clients with IDs [32, …, 50]), the merging weights evolve more slowly, even after 500 epochs. This is because the largest cluster has the highest number of clients (19 clients) and requires more global models, resulting in a more uniform merging weight distribution. This does not mean that large clusters are difficult to converge. It simply indicates that the weight updates are more evenly distributed within larger clusters. These small weight changes within large clusters are sufficient to significantly distinguish large clusters from other clusters.
@@
Appendix B More Experiments on FL with Foundation Models
More Discussion on Personalized LLM Datasets
Dataset for Persona-Aligned Instruction Following. There exist several baselines (Ye et al. 2024b; Kuang et al. 2024; Ye et al. 2024a) for federated foundation models. However, these models are primarily designed for instruction fine-tuning, and their goal is to learn a global model to follow general instructions rather than personalize the LLM’s behavior to each client. The evaluation metrics in these baselines are often datasets or evaluation models independent of the FL training environment. Therefore, to develop personalized federated LLMs, it is essential to focus on datasets that reflect individual characteristics—such as culture (Li et al. 2024; Shi et al. 2024), style (Fisher et al. 2024; Hu et al. 2022), and personality traits (Wang et al. 2023; Shao et al. 2023). In this paper, we use RoleBench (Wang et al. 2023), a collection of $100$ characters. The role-specific instructions and responses are used as the personalized dataset for each client. Each client is assigned one specific character, and we randomly sample $50$ characters out of $100$, resulting in $50$ clients. For each selected character, we randomly select half of the original training dataset and test dataset as the client-specific dataset. Table 5 presents some input-output examples. Following the evaluation metric in (Wang et al. 2023), ROUGE-L is used as the evaluation metric. At each communication round, $20$ clients are randomly sampled out of $50$.
Task Dataset
Paraphrase glue_qqp
Entailment snli
Structure to text web_nlg_en
Text formatting fix_punct
Linguistic acceptability cola
Word disambiguation wic
Coreference definite_pronoun_resolution
Question classification trec
Table 4: Federated Flan dataset for personalized instruction Tuning.
Dataset for Task-Heterogeneous Instruction Following. We construct a federated dataset from Flan (Wei et al. 2021), which is a collection of various classic NLP tasks from over 60 datasets for instruction tuning. Following (Long et al. 2024), we randomly select $8$ NLP tasks, one task for one client, resulting in $8$ total clients. For each task, we randomly select $300$ training samples and $200$ test samples from the original dataset. ROUGE-1 is used as the evaluation metric. We perform full client selection.
Role Jack Sparrow — A witty and eccentric pirate captain from the Pirates of the Caribbean film series.
Question How did you manage to escape from the casket?
Answer Ah, well, you see, I always have a few tricks up my sleeve, or in this case, up my coffin. It wasn’t easy, though, as I had to wait for just the right moment to break open the casket without drawing too much attention to myself. But, as they say, desperate times call for desperate measures, and I’ve always been good at thinking on my feet.
Role Stephen Hawking — A renowned theoretical physicist known for his work on black holes and cosmology.
Question What motivated you to pursue a career in cosmology?
Answer I’ve always been fascinated by the mysteries of the universe, and I wanted to dedicate my life to uncovering its secrets. I believe that through science, we can better understand our place in the world and how everything is interconnected.
Table 5: Roleplay examples in the federated RoleBench dataset, showing character-specific questions and responses. Each question is an input prompt to an LLM, and the corresponding answer is the generated response.
Baseline Settings for Foundation-model FL Experiments
pFedMoE (Yi et al. 2024) employs a personalized router to balance contributions from the global and local models. Following this design, we introduce a router for each pair of global and local LoRA modules. Each router is implemented as a linear layer followed by a softmax activation. The router takes as input the same embedding used by the LoRA modules and outputs a routing score. This score is then used to compute a weighted average of the outputs from the global and local LoRA modules, yielding the final LoRA output.
Table 6: Parameter scale settings for LoRA modules on the Federated RoleBench dataset. Parameter scale is measured relative to Rank-8.
Multi-Model Description Rank-8×5 Rank-8×10 Rank-8×15 Rank-8×20 Rank-8×25 Rank-8×30
Parameter Scale 5 10 15 20 25 30
Single-Model Description Rank-8 Rank-16 Rank-24 Rank-32
Parameter Scale 1 2 3 4
FedEM (Marfoq et al. 2021) requires a personalized router for a group of global models. Similar to pFedMoE, we add a router for each set of global LoRAs, and the output dimension of the router is the number of global LoRAs. The input embedding of the router is also the input to the LoRA module. The router produces routing scores, which are then used to compute a weighted combination of the global LoRA outputs, yielding the final LoRA output.
Figure 8: Server parameter usage vs. accuracy on the federated RoleBench dataset. The values on the x-axis are shown as multiples of the parameter count of a Rank-8 LoRA. The highest and second-highest accuracy points are marked in black and gray, respectively.
FedIT-FFALoRA As in (Sun et al. 2024), we only transmit and optimize the $B$ matrix of LoRA. The $A$ matrix is initialized and frozen on each client.
FedMerge Due to the varying lengths of predicted texts across clients, the gradient magnitudes returned to the server differ significantly. As a result, the merging weights are updated at uneven rates, with some parameters receiving large updates while others remain almost unchanged. To address this non-IID issue, we normalize the gradient of each client’s corresponding merging weight, ensuring that all clients contribute gradients with consistent magnitudes while preserving their unique update directions. We set the learning rate to 0.01, consistent with the learning rate used in classical FL settings.
Results for Federated RoleBench Dataset
Following Table 6, we vary Single-Model methods by LoRA rank and Multi-Model methods by the number of LoRA modules. FedEM has to optimize all global LoRA modules simultaneously. Due to our limited computational resources, we only vary the number of shared LoRA modules for FedEM from $2$ to $10$. Since there are $50$ clients, we compute the average accuracy over the bottom-half clients to evaluate each method’s ability to handle the worst-performing characteristics. The server parameter usage vs. accuracy for the federated RoleBench dataset is shown in Fig. 8. FedMerge outperforms baselines, demonstrating its effectiveness in supporting the worst-performing clients.